OpenStudy (christian_10_):

What are the square roots of 576? A. −14 and 14 B. −24 and 24 C. −28 and 28 D. −288 and 288

Mehek (mehek14):

@Christian_10_ which one do you think?

OpenStudy (christian_10_):

I'm not exactly sure on this one!

OpenStudy (er.mohd.amir):

x*x=x^2 multiply all ans get result one way to solve also other ways also

Mehek (mehek14):

what's \(14^2\) or \(-14^2\) = ?

OpenStudy (lanhikari22):

Can't you check from the answers? You know that a number is a square root of another when the number multiplied by itself gives the other number. 25 has a square root of 5, or -5. 5 * 5 = ? -5 * -5 = ?

OpenStudy (mathmale):

@Mehek14 : I need to comment on your post that includes -14^2 . It's mandatory to enclose that -14 in parentheses; otherwise your result will have the wrong sign.